The sum of the first and second term of an AP is 4,if the tenth term is 19 find the sum of the fifth and sixth terms
Answer by richwmiller(17219)

an = a + (n - 1)d
4= 2a + d,
19= a + 18d
a = 53/35, d = 34/35
check
53/35+53/35+34/35
140/35=4
53/35+612/35=665/35
s = 53/35 + (4)34/35+ 53/35 + (5)34/35
s=412/35 =11 27/35